OpenStudy (anonymous):

4x2 + 4x + 1 = (2x + 1)2

OpenStudy (anonymous):

true of false

mathslover (mathslover):

Is it : \[\large{4x^2 + 4x + 1 = (2x+1)^2}\] or \[\large{4x^2 + 4x + 1 = (2x+1)2}\] ?

OpenStudy (anonymous):

the 1st one

mathslover (mathslover):

OK so do you know the formula for \(\large{(a+b)^2}\) ?

OpenStudy (anonymous):

i think so

mathslover (mathslover):

What is it ?

OpenStudy (anonymous):

ummmm

OpenStudy (anonymous):

this is hard

mathslover (mathslover):

@Mertsj will continue to help you as I have got urgent work Thanks a lot Mertsj

OpenStudy (mertsj):

yw

OpenStudy (mertsj):

(a+b)^2 = a^2+2ab+b^2

OpenStudy (mertsj):

We suspect our trinomial is of this form since the first term is a perfect square and the last term is a perfect square.

OpenStudy (mertsj):

So then we think: if 4x^2 is a squared, then a is 2x. If b^2 is 1 then b is 1 and 2ab is 2(2x)(1) which is the middle term.

OpenStudy (mertsj):

So we know we have a trinomial square and the factors are (2x+1)^2